Find the 7th term of an A.P is 15 and the fourth term is 9. Find the sequence of first term

1
Expert's answer
2021-11-16T06:08:28-0500

In an arithmetic progression the formula is

 = a1 + ( n - 1 ) d


where


a1= the initial term


d = the common difference of successive members is d,


an = the nth term


In this case:

a4 = a1 + ( 4 - 1 ) d = 15


a1 + 3 d = 15


a7 = a1 + ( 7 - 1 ) d = 27


a1 + 6 d = 27


Now you must solve system:


a1 + 3 d = 15


a1 + 6 d = 27

solving simultaneously gives

a1=3

d= 4

The AP is

3,7,11,15,19,23, 27

The sequence of first term

= a1 + ( n - 1 ) d

= 3 + (1-1)4

where n= first term

=3